OpenStudy (ribhu):

\[\int\limits_{0}^{oo}logx/ 1+x^{2}dx\] will FAN and Medal

OpenStudy (ribhu):

@Jhannybean please solve

OpenStudy (ribhu):

@Gabylovesyou

OpenStudy (jhannybean):

\[\int_0^\infty \frac{\log(x)}{1+x^2}dx\]This?

OpenStudy (jhannybean):

can this be done by parts...

OpenStudy (jhannybean):

Or LH rule?

OpenStudy (ribhu):

if u solve this.

OpenStudy (ribhu):

@Jhannybean still not able to get

OpenStudy (freckles):

Are you just suppose to determine if it is divergent or convergent ?

OpenStudy (freckles):

\[x^2+1>x^2 \\ \frac{1}{x^2+1}<\frac{1}{x^2} \\ \frac{\ln(x)}{x^2+1}<\frac{\ln(x)}{x^2}\] if so try comparision test

OpenStudy (ribhu):

@Preetha ma'am please help!

OpenStudy (ribhu):

can this question be solved?? @dan815 @nincompoop @Kainui please help me in this.

OpenStudy (ribhu):

i have been trying everything but couldn't solve it

ganeshie8 (ganeshie8):

substitute \(x = \tan u \implies \dfrac{dx}{1+x^2} = du\), the integral becomes : \[\int_0^\infty \frac{\ln(x)}{1+x^2}dx = \int\limits_{0}^{\pi/2} \ln(\tan u) \]

OpenStudy (ribhu):

this comes out to be zero.

ganeshie8 (ganeshie8):

Next notice below to conclude the integral evaluates to 0 : \[ \int\limits_{0}^{\pi/2} \ln(\tan u) du= \int\limits_{0}^{\pi/2} \ln(\tan (\pi/2-u))du =\int\limits_{0}^{\pi/2} \ln(\cot u)du = -\int\limits_{0}^{\pi/2} \ln(\tan u) du \]

ganeshie8 (ganeshie8):

\(I = -I \) a thing is equal to its negaitve thats only possible when it is 0

OpenStudy (ribhu):

if this is indefinite integral?

ganeshie8 (ganeshie8):

then i think we simply say it is not possible to represent in elementary functions :P http://www.wolframalpha.com/input/?i=%5Cint++%5Cfrac%7B%5Clog%28x%29%7D%7B1%2Bx%5E2%7Ddx

OpenStudy (ribhu):

ok.